youngsponge007:

How far is the airplane from the runway? Round to the nearest tenth. 10,000 32,000

youngsponge007:

same concept as the last one

Tonycoolkid21:

ok got it

Tonycoolkid21:

@youngsponge007

youngsponge007:

sooooo Whats the awnser

Tonycoolkid21:

did you see the picture

youngsponge007:

yeah but ion see it

Tonycoolkid21:

i wil post it again

youngsponge007:

ok

youngsponge007:

i see the pic but the awser not there

Tonycoolkid21:

k hold on i will give it to you

youngsponge007:

ok

youngsponge007:

no its not

snowflake0531:

We know that they pythagorean theorem is \[a^2 + b^2 = c^2\] while a is 10,000, and b is 32,000 So, plug a and b in, and tell me what you get

youngsponge007:

no its not

youngsponge007:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 We know that they pythagorean theorem is \[a^2 + b^2 = c^2\] while a is 10,000, and b is 32,000 So, plug a and b in, and tell me what you get \(\color{#0cbb34}{\text{End of Quote}}\) so 20,000 and 64,000

snowflake0531:

No \[10000^2 + 32000^2 = c^2\] What does the left equal

youngsponge007:

5000

snowflake0531:

do you know exponents

youngsponge007:

nope

youngsponge007:

i keep getting 33526.11

snowflake0531:

what grade are you in

youngsponge007:

i know them but not well

snowflake0531:

10000^2 = 10,000* 10,000 32,000^2 = 32,000 * 32,000

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @youngsponge007 i keep getting 33526.11 \(\color{#0cbb34}{\text{End of Quote}}\) correct

snowflake0531:

now round to the nearest tenths

youngsponge007:

oh ok

youngsponge007:

so just 33526

snowflake0531:

No, that would be without decimal you still need the tenths place

youngsponge007:

ok so 33526.1

snowflake0531:

yes

Tariyah:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 yes \(\color{#0cbb34}{\text{End of Quote}}\) who u sayin yes to

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @youngsponge007 ok so 33526.1 \(\color{#0cbb34}{\text{End of Quote}}\) yes

Tonycoolkid21:

ok, i got the answer so to do this you need to use Pythagoreans theorem This says a^2+b^2=c^2 plug in the two numbers for a and b (32000)^2+(10000)^2=C^2 1124000000=C^2 then find the square root c=+-33526.109. since it is the distance the answer can only be positive Answer: 33526.109ft

Tonycoolkid21:

@youngsponge007 anymore questions

youngsponge007:

nope

Tonycoolkid21:

k

youngsponge007:

no u didn't lol

Tonycoolkid21:

but you did not explain it you cant just give people answer with out explaining it

Tonycoolkid21:

@youngsponge007 also thank you for the medal :)

youngsponge007:

yw